# FaceEditTalker: Interactive Talking Head Generation with Facial Attribute Editing

### v1

@@ Abstract
Recent advances in audio-driven talking head generation have achieved impressive results in lip synchronization and emotional expression. However, they largely overlook the crucial task of facial attribute editing. This capability is crucial for achieving deep personalization and expanding the range of practical applications, including user-tailored digital avatars, engaging online education content, and brand-specific digital customer service. In these key domains, the flexible adjustment of visual attributes—such as hairstyle, accessories, and subtle facial features—is essential for aligning with user preferences, reflecting diverse brand identities, and adapting to varying contextual demands. In this paper, we present FaceEditTalker, a unified framework that enables controllable facial attribute manipulation while generating high-quality, audio-synchronized talking head videos. Our method consists of two key components: an image feature space editing module, which extracts semantic and detail features and allows flexible control over attributes like expression, hairstyle, and accessories; and an audio-driven video generation module, which fuses these edited features with audio-guided facial landmarks to drive a diffusion-based generator. This design ensures temporal coherence, visual fidelity, and identity preservation across frames. Extensive experiments on public datasets demonstrate that our method outperforms state-of-the-art approaches in lip-sync accuracy, video quality, and attribute controllability. Project page: https://peterfanfan.github.io/FaceEditTalker/
1 Introduction
@@
In recent years, audio-driven talking head generation  [38, 35, 53, 10, 16, 45, 30] has made remarkable progress and found widespread applications in domains such as virtual reality[28, 20], animation production[25, 51], online education[20], digital humans[16], and film post-production[28]. These technologies enable virtual characters to exhibit more natural and realistic speaking behaviors by synchronizing facial movements with audio input. However, most existing approaches primarily focus on lip synchronization [38, 35, 53, 10] and emotional expression [14, 50, 32, 13], while largely overlooking the important functionality of facial attribute editing.
@@
Facial attribute editing is essential for audio-driven video generation due to its strong practical relevance. Beyond accurate audio-visual synchronization, users often require flexible control over visual appearance, including expressions, hairstyles, age, gender, and accessories like glasses. For example, virtual idols may need to adapt to different audience preferences, and digital customer service agents may need to reflect distinct brand identities. Dynamic and fine-grained attribute control can greatly enhance personalization and user engagement.
Although image-level facial attribute editing methods such as GAN-based or text-driven approaches [21, 23] have achieved initial success in static image generation tasks, extending these methods to video generation, particularly in talking-head videos, remains challenging. The main obstacles include maintaining temporal consistency of facial attributes, where edits must not only preserve realism within individual frames but also ensure smooth transitions across frames to avoid visual flickering or discontinuities, which is essential for producing natural video sequences [56, 49, 24]. Moreover, preserving audio-driven facial dynamics during attribute manipulation is crucial, as any inconsistency can disrupt accurate lip synchronization and natural facial motion.
To solve these problems, we propose FaceEditTalker, a novel method that seamlessly integrates audio-driven facial animation with controllable facial attribute editing. Our approach enables users to flexibly adjust facial attributes during video generation, while preserving high image fidelity, smooth facial dynamics, and accurate audio-visual alignment.
@@
(1) We propose a novel framework named FaceEditTalker, which seamlessly unifies facial attribute editing and audio-driven talking head generation. This framework enables fine-grained manipulation of attributes such as hair feature, facial structure, and accessories, while maintaining high-quality lip synchronization and natural motion dynamics.
@@
(2) We propose an innovative two-stage heterogeneous latent diffusion model to address challenges in editability and consistency. Through editable latent feature extraction and feature re-injection, it enables highly flexible zero-shot editing while preserving identity integrity and natural dynamics.
@@
(3) We conduct extensive experiments on multiple public datasets such as HDTF[60] and VoxCeleb2[33], demonstrating that our method significantly outperforms state-of-the-art baselines in terms of video quality, lip synchronization accuracy, keypoint alignment error, and identity consistency.
2 Related Work
@@
Audio-driven Talking Head Generation. Recent methods for audio-driven talking head generation have made remarkable progress, emphasizing realism, identity preservation, and expression diversity. Early approaches [38, 35, 53, 10] primarily adopt encoder-decoder architectures to map audio signals to lip movements. While effective to some extent, these methods often suffer from blurred textures and weak identity preservation due to limited fusion strategies. To enhance realism, NeRF-based methods [16, 45, 30, 31, 37] and 3D Gaussian Splatting methods [2, 57, 29, 15] both model 3D geometry for more lifelike appearances; however, they typically require long video sequences and come with high computational costs. Another line of work leverages facial landmarks or 3D priors [61, 7, 59, 58] to disentangle speech content from identity features, improving controllability but often sacrificing fine-grained details in regions like lips and teeth. Recently, diffusion-based models [46, 48, 54, 44, 41, 6, 55, 17, 18, 2] have emerged as powerful solutions for high-quality and diverse talking head generation, with many of these approaches also leveraging facial landmarks as driving signals to enhance controllability and expression alignment. In our method, facial landmarks are adopted as controllable priors to guide a diffusion-based generator, achieving precise lip synchronization and identity preservation while supporting flexible facial attribute editing without compromising speech-driven facial dynamics.
Facial Attribute Editing. Facial attribute editing aims to manipulate specific attributes (e.g., age, hairstyle, glasses) while preserving identity. StyleGAN-based methods [21, 23, 43, 1] leverage latent space disentanglement to enable controllable editing, while CLIP-guided approaches [42, 36] introduce semantic alignment between language and image, allowing intuitive text-driven attribute modifications. Diffusion models offer enhanced control and fewer artifacts for editing tasks [5, 11, 40], but extending them to videos introduces temporal consistency challenges due to frame-wise randomness. To address this, methods like Latent-Transformer [56], STIT [49], and Diffusion-Video-Autoencoders [24] decompose identity and motion or apply global-local refinements to preserve coherence across frames. In our method, we perform linear transformations in the diffusion latent space while keeping the first frame fixed, ensuring both high-quality attribute editing and consistent facial dynamics across video frames.
3 Method
@@
Figure 2: Overview of the inference process of our proposed framework FaceEditTalker. The framework consists of two main modules: (a) Image Feature Space Editing Module, which extracts editable semantic and stochastic codes from the reference image using a dual-layer latent encoding structure. Fine-grained attribute manipulation is enabled through optional spatial editing on the semantic codes. (b) Audio-Driven Video Generation Module, which leverages the audio input to infer driving landmarks. During the diffusion process, the stochastic codes guide dynamic generation, while the semantic codes serve as conditional inputs to ensure attribute consistency and visual fidelity throughout the video. The training procedure is detailed in Section 3.5 and Appendix A.
3.1 Overview
@@
Our proposed framework, FaceEditTalker, introduces an innovative two-stage heterogeneous latent diffusion model to address challenges in editability and consistency. It consists of two tightly coupled modules: the Image Feature Space Editing Module and the Audio-Driven Video Generation Module. The Image Feature Space Editing Module extracts editable semantic and stochastic codes from the reference image using a dual-layer latent encoding structure, enabling fine-grained control over facial attributes, which can be further edited using text-guided linear classifiers. These features are then passed to the Audio-Driven Video Generation Module, where synchronized audio-driven landmarks guide a diffusion-based generative process to produce high-quality, temporally coherent talking head videos with consistent identity and natural lip-sync.
3.2 Task Formulation

### v2

# {textblock*}

@@ Abstract
Recent advances in audio-driven talking head generation have achieved impressive results in lip synchronization and emotional expression. However, they largely overlook the crucial task of facial attribute editing. This capability is indispensable for achieving deep personalization and expanding the range of practical applications, including user-tailored digital avatars, engaging online education content, and brand-specific digital customer service. In these key domains, flexible adjustment of visual attributes, such as hairstyle, accessories, and subtle facial features, is essential for aligning with user preferences, reflecting diverse brand identities and adapting to varying contextual demands. In this paper, we present FaceEditTalker, a unified framework that enables controllable facial attribute manipulation while generating high-quality, audio-synchronized talking head videos. Our method consists of two key components: an image feature space editing module, which extracts semantic and detail features and allows flexible control over attributes like expression, hairstyle, and accessories; and an audio-driven video generation module, which fuses these edited features with audio-guided facial landmarks to drive a diffusion-based generator. This design ensures temporal coherence, visual fidelity, and identity preservation across frames. Extensive experiments on public datasets demonstrate that our method achieves comparable or superior performance to representative baseline methods in lip-sync accuracy, video quality, and attribute controllability. Project page: https://peterfanfan.github.io/FaceEditTalker/. We will release the source code to the public upon acceptance.
Index Terms:
audio-driven talking head generation, facial landmark, semantic feature disentanglement, facial attribute editing.
I Introduction
@@
In recent years, audio-driven talking head generation [1, 2, 3, 4, 5, 6] has achieved remarkable progress and found widespread applications in domains such as virtual reality [7, 8], digital humans [4], online education [8], animation production [9], and film post-production [7]. These methods enable virtual characters to synchronize facial movements with audio input, producing natural speaking behaviors. However, most existing approaches primarily focus on lip synchronization [1, 2, 3, 10, 11] and emotional expression [12, 13, 14, 15, 16, 17], while largely overlooking the important functionality of controllable facial attribute editing.
@@
Facial attribute editing is essential for audio-driven video generation due to its strong practical relevance. Beyond accurate audio-visual synchronization, users often require precise and flexible control over visual appearance, including expressions, hairstyles, age, gender, makeup, and accessories like glasses. For example, virtual idols may need to adapt to different audience preferences, and digital customer service agents may need to reflect their distinct brand identities. Dynamic and fine-grained attribute control can greatly enhance user engagement and personalization.
Previous research on facial attribute editing has been extensive, initially focusing on static face images. GAN-based methods have achieved significant success in this domain, with representative examples including StyleGANs [18, 19], which leverage a highly disentangled latent space to enable realistic and controllable facial edits. Naturally, researchers have attempted to extend these techniques to video generation, which introduces new challenges in maintaining facial detail, temporal consistency, and overall video quality. (1) Poor facial detail: Although GAN-based methods employ strategies such as frame alignment and fine-tuning to preserve temporal coherence during frame-by-frame editing, these approaches can still result in misalignment artifacts and inconsistent facial details [20], as well as background flicker [21] and other visible video artifacts. (2) Temporal discontinuity in editing: Existing methods often suffer from temporal artifacts, such as visual flickering [22] or fluctuations of dynamic attributes (e.g., beard, eyeglasses) during motion [23, 24]. Variations in head pose can further compromise temporal consistency, resulting in non-smooth or perceptually unstable edits. These limitations are largely attributable to the intrinsic capacity constraints of GAN-based models, which hinder their ability to accurately encode and transfer the complex information embedded in both source and target attribute frames. Although diffusion-based approaches generally yield higher video fidelity and more robust attribute manipulation than GAN-based methods, they are nevertheless susceptible to both imperfect facial detail [25] and temporal inconsistency [26, 27, 25].
To address these limitations, we propose FaceEditTalker, a novel framework combining audio-driven talking head generation with controllable facial attribute editing. We adopt a dual-layer latent encoding structure [28] to jointly model high-level semantics and low-level textures, where the semantic encoder conditions the reference image to guide DDIM for accurate facial reconstruction. A linear classifier is trained on the attribute semantic code to produce an attribute vector stored in a label-vector library; fusing this vector with the original semantic code yields an edited semantic code that guides DDIM to generate the desired attributes. Unlike StyleGAN-based methods [20, 24], which often sacrifice reconstruction quality for precise editing, our approach achieves near-perfect face reconstruction while preserving fine-grained details. Furthermore, we adopt the landmark predictor from the representative method [29] to accurately infer landmark features, enabling joint guidance by semantic code and landmarks to prevent facial jitter and attribute fluctuations, ensuring temporal consistency. Our framework seamlessly integrates these components to achieve high-fidelity and editable talking head generation. Extensive experiments on multiple public datasets demonstrate superior performance in video quality, keypoint alignment, and identity preservation.
@@
  • We propose FaceEditTalker, the first framework that seamlessly unifies facial attribute editing and audio-driven talking head generation, enabling fine-grained manipulation of attributes such as hair, facial structure, and accessories, while maintaining natural lip movements and facial dynamics.
@@
  • We introduce a novel two-stage heterogeneous latent diffusion model to address the challenges of editing capability and consistency, enabling highly flexible zero-shot editing while effectively preserving identity integrity and temporal coherence.
@@
  • We conduct extensive evaluations on multiple public datasets, demonstrating that our method outperforms existing baselines in video quality, lip synchronization, keypoint alignment, and identity preservation.
II Related Work
II-A Audio-driven Talking Head Generation.
@@ II Related Work
Recent methods for audio-driven talking head generation have made remarkable progress, emphasizing realism, identity preservation, and expression diversity. Early approaches [1, 2, 3] primarily adopt encoder-decoder architectures to map audio signals to lip movements. Although effective to some extent, these methods often suffer from blurred textures and weak identity preservation due to limited fusion strategies. To enhance realism, NeRF-based methods [4, 5, 6, 30, 31] and 3D Gaussian Splatting methods [32, 33, 34, 35, 36, 37] model 3D geometry for more lifelike appearances; however, they typically require long video sequences and come with high computational costs, limiting their practicality in real-time scenarios. Another line of work leverages facial landmarks or 3D priors [38, 39, 40, 41, 42] to disentangle speech content from identity features, improving controllability but often sacrificing fine-grained details in critical regions such as lips and teeth. More recently, diffusion-based models [32, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] have emerged as a promising direction for high-quality and expressive talking head generation. Unlike GAN-based approaches that generate frames in a single forward pass, diffusion models iteratively denoise random noise under the guidance of conditioning signals such as audio and landmarks, enabling more precise control over motion dynamics and temporal consistency. Operating in structured control pipelines, these models are capable of synthesizing realistic lip movements, nuanced expressions, and coherent head pose.
In our method, facial landmarks are adopted as controllable priors to guide a diffusion-based generator, enabling precise lip synchronization and identity preservation while supporting flexible facial attribute editing without compromising speech-driven facial dynamics. Furthermore, the semantic code encoded from the reference image is incorporated as an additional control condition to guide the generation process, ensuring the preservation of fine-grained facial details.
@@
Figure 2: Overview of the inference process of our proposed framework FaceEditTalker. The framework consists of two main modules: (a) Image Feature Space Editing Module, which extracts editable semantic and stochastic codes from the reference image using a dual-layer latent encoding structure. Fine-grained attribute manipulation is enabled through optional spatial editing on the semantic codes. (b) Audio-Driven Video Generation Module, which leverages the audio input to infer driving landmarks. During the diffusion process, the stochastic codes guide dynamic generation, while the semantic codes serve as conditional inputs to ensure attribute consistency and visual fidelity throughout the video. The training procedure is detailed in Section III-E and Section IV-B.
II-B Facial Attribute Editing.
Facial attribute editing focuses on modifying specific facial characteristics, such as age, hairstyle, and glasses, while preserving the subject’s identity. StyleGAN-based methods [18, 19, 56, 57] achieve controllable editing through latent space disentanglement, while CLIP-guided approaches [58, 59] introduce semantic alignment between text and images, enabling intuitive language-driven modifications. The emergence of diffusion models has further expanded the possibilities for producing realistic and expressive facial animations, offering enhanced editing control and fewer visual artifacts [28, 60, 61]; however, extending these models to video sequences introduces challenges in maintaining temporal consistency due to frame-wise stochasticity, which can cause attribute variations across frames and lead to inconsistency. To address this, Latent Transformer [23] performs optical flow alignment and refines latent codes in StyleGAN $\mathcal{W}^{+}$ with identity- and attribute-preserving regularization to achieve temporally stable frame-wise editing. STIT [24] encodes adjacent frames into smoothly varying latent codes, enforces global identity via PTI, and performs stitching-based refinement to integrate edits without relying on explicit temporal loss. Diffusion Video Autoencoders [26] represent a video using a shared identity vector and per-frame motion/background vectors; editing the shared identity ensures temporal coherence and consistent facial attributes across frames.
Despite these advances, achieving temporally consistent and high-fidelity facial attribute editing remains a key challenge in video-based editing. To address this, we introduce a motion module within the DDIM framework that employs Temporal Self-Attention across multiple resolutions with temporal position encoding to capture frame-to-frame dependencies. Attribute editing is performed in two stages: attribute vectors are first derived using a trained linear classifier, then used to adjust the semantic code toward the target attribute. The modified semantic code subsequently guides the DDIM to generate videos that retain the first-frame identity while exhibiting the desired attribute. This design enables precise, high-fidelity edits with coherent facial dynamics across frames.
III Method
In this section, we provide a comprehensive and detailed description of the FaceEditTalker framework. Section III-A presents an overview of the overall architecture. Section III-B formulates the task and outlines the workflow of the method. Section III-C details the image feature space editing module for facial attribute manipulation, while Section III-D describes the audio-driven video generation module. Finally, Section III-E clarifies the training objectives, loss functions, and the inference procedure.
Figure 3: Workflow of the Facial Editing Process. The attribute direction vector $w_{att}$ is learned via a linear classifier. The original semantic code $z_{sem}$ is then linearly transformed along this direction with a strength factor $\alpha$ to produce the edited semantic code containing the desired attribute.
III-A Overview
@@
Our proposed framework, FaceEditTalker, consists of two tightly coupled modules: the Image Feature Space Editing Module and the Audio-Driven Video Generation Module as shown in Fig. 2. The Image Feature Space Editing Module extracts editable semantic and stochastic codes from the reference image using a dual-layer latent encoding structure, enabling fine-grained control over facial attributes, which can be further edited using text-guided linear classifiers. These features are then passed to the Audio-Driven Video Generation Module, where synchronized audio-driven landmarks guide a diffusion-based generative process to produce high-quality, temporally coherent talking head videos with consistent identity and natural lip-sync.
III-B Task Formulation
We first introduce the notations used in our formulation. Let $x_{ref}\in\mathbb{R}^{H\times W\times 3}$ denote the reference image of the target person, $A^{1:T}=(a^{1},\dots,a^{T})$ the extracted audio features, $y$ the facial attribute labels, $z_{\text{att}}$ the attribute code and $z_{\text{sem}}$ the semantic code. The driving landmark sequence is represented as $L^{1:T}=(l^{1},l^{2},\dots,l^{T})\in\mathbb{R}^{T\times H\times W\times 3}$, and the generated video as $\hat{V}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T}\}\in\mathbb{R}^{T\times H\times W\times 3}$.
For facial attribute editing, we leverage a labeled face dataset. Each image is encoded by a semantic encoder into an attribute vector $z_{\text{att}}$, which, together with its label $y$, is used to train a linear classifier $C$ that learns the mapping between labels and attribute vectors. During editing, $y$ serves as a key to retrieve the corresponding vector, and the semantic representation $z_{\text{sem}}$ is modified with a manipulation strength $\alpha$, formally expressed as
$z_{\text{sem}}^{\prime}=C(z_{\text{sem}},y,\alpha).$ (1)
For audio-driven talking head generation, we adopt the pre-trained wav2vec model [62] to extract audio features $A^{1:T}$, which are processed by a multiscale landmark prediction network to generate the driving landmark sequence $L^{1:T}$. Given $x_{ref}$ and $L^{1:T}$, the objective is to synthesize a realistic talking-head video $\hat{V}$ that preserves the identity of $x_{ref}$ while following the motion dynamics of $L^{1:T}$. This process is formulated as
$\hat{V}=g(x_{ref},L^{1:T},z_{\text{sem}}^{\prime}),$ (2)
where $g$ denotes the proposed generative model. Details of each component are presented in the following sections.
III-C Image Feature Space Editing Module
The semantic encoder $E_{\text{sem}}$ extracts global facial semantics from the input image, encoding them into low-dimensional vectors akin to StyleGAN’s style vectors, enabling linear transformations for attribute editing. As shown in Fig. 3, to enable text-driven editing and better align with the illustrated process, we employ the semantic encoder to extract semantic codes from images annotated with target attributes, thus constructing an attribute library. In the next step, the text input is processed in conjunction with the attribute library by a linear classifier to obtain the attribute direction vector $w_{\text{att}}$. The linear classifier is implemented as a single-layer MLP, and the attribute direction vector corresponds to its weight parameters. Subsequently, the original semantic code $z_{\text{sem}}$ is transformed into the dimension as $w_{\text{att}}$, enabling $w_{\text{att}}$ to perform a linear transformation that steers the semantic code towards the desired attribute. Given a semantic code $z_{\text{sem}}$ and an attribute direction vector $w_{\text{att}}$, attribute manipulation is expressed as:
$z^{\prime}_{\text{sem}}=z_{\text{sem}}+\alpha\cdot w_{\text{att}},$ (3)
where $\alpha$ controls the intensity of the change, and $z^{\prime}_{\text{sem}}$ denotes the edited semantic code.
In this attribute editing process, any of the 40 predefined attributes can be flexibly selected. By specifying the desired attribute at the corresponding code position, the editing process described above can be applied to achieve attribute manipulation on the image. Furthermore, it is also possible to design automated scripts to systematically generate edited results for all attributes. To ensure consistency in non-target areas, we use the reconstruction loss $L_{\text{rec}}$ to measure the difference between the edited latent code $z^{\prime}_{\text{sem}}$ and the static code $z_{\text{static}}$ representing unedited regions. The reconstruction loss $L_{\text{rec}}$ is defined as:
$L_{\text{rec}}=|z^{\prime}_{\text{sem}}-z_{\text{static}}|.$ (4)
The stochastic encoder, designed as a Unet-based diffusion model [63]. First, the image encoder $E_{img}$ encode the input image into an initial latent representation $Z_{0}$. This representation undergoes a noise diffusion process to capture fine details, resulting in the final latent representation $Z_{T}$. The parameter $\alpha_{t}$ controls the proportion of the original signal retained in each diffusion step, regulating the balance between signal preservation and noise injection during the forward process. We refer to this complete encoding and diffusion procedure as a stochastic encoder. The forward process is defined as:
$z_{t+1}=\sqrt{\alpha_{t+1}}z_{t}+\sqrt{1-\alpha_{t+1}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (5)
During the forward diffusion process, the initial latent representation $Z_{0}$, obtained from the input image, is progressively noised using our defined stochastic encoder. This encoder is specifically designed to implement the forward diffusion process, in which Gaussian noise is systematically added to $Z_{0}$ over a sequence of time steps according to a predefined noise schedule. As a result, the representation gradually transitions into a highly stochastic latent code $Z_{T}$, which captures rich local detail variations essential for realistic facial generation in the subsequent reverse process.
III-D Audio-Driven Video Generation Module
As shown in Fig. 2, this module generates video using audio input $S_{\text{audio}}$, frame sequence $V=\{x_{1},x_{2},\dots,x_{T}\}$, semantic code $z_{\text{sem}}$ (irrespective of whether it is edited) and stochastic code $Z_{T}$. To perform audio-to-landmark prediction, we employ two networks: a landmark predictor $P_{\text{ldm}}$ and a landmark feature extractor $E_{\text{ldm}}$.
First, the audio input $S_{\text{audio}}$ is first processed by a pre-trained Wav2Vec model to extract audio features $A^{1:T}$:
$A^{1:T}=Wav2Vec(S_{audio}).$ (6)
Landmark Predictor: Inspired by AniPortrait [29], which demonstrates strong lip-sync accuracy and temporal consistency, we define an audio-driven landmark predictor $P_{\text{ldm}}$ to generate a facial landmark sequence $L^{1:T}$ from the audio features $A^{1:T}$ and the reference image $x_{\text{ref}}$:
$L^{1:T}=P_{\text{ldm}}(A^{1:T},x_{\text{ref}}).$ (7)
Landmark Feature Extractor: The facial landmark sequence $L^{1:T}$ is processed by a landmark feature extractor to obtain the corresponding landmark features $F^{1:T}$:
$F^{1:T}=E_{ldm}(L^{1:T},x_{ref}),$ (8)
where $E_{\text{ldm}}$ employs multiscale strategies and cross-attention mechanisms to fuse the landmark sequence $L^{1:T}$ with the reference landmark sequence $l_{\text{ref}}$ extracted from the reference image $x_{\text{ref}}$, producing landmark features $F^{1:T}$. This design enables the model to capture multilevel facial dynamics and strengthen feature correlations, thus improving precision and temporal consistency in video generation, as further validated by the ablation results reported in Table V.
Subsequently, the facial landmark features $F^{1:T}$ are fused with the stochastic code $Z_{T}$ through a residual connection, producing the feature representation $K$, which is then used as the input to the diffusion model:
$K=Concat(F^{1:T},Z_{T})$ (9)
During the diffusion model sampling process, we employ conditional DDIM [64], where high-level semantic information $z_{\text{sem}}$ is incorporated as a global facial attribute control signal, while the feature $K$ provides dynamic motion information. This design ensures stable expression of semantic attributes alongside synchronized audio-driven facial movements. The DDIM reverse sampling at timestep $t$ is formulated as:
$z_{t-1}=\sqrt{\alpha_{t}}z_{t}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}(z_{t},z_{\text{sem}},F,t),$ (10)
where $z_{t}$ represents the noisy data in the timestep $t$, and $\epsilon_{\theta}$ denotes the conditional denoising network that guides the denoising process based on the latent variable $z_{\text{sem}}$ and the input feature $K$. The conditional feature injection in the diffusion model employs a dual mechanism:
  • Semantic code $z_{\text{sem}}$: injected via cross-attention layers to provide global facial attribute control.
  • Input feature $K$: fused from the inputs to deliver local motion control.
Furthermore, multiscale feature fusion ensures that control signals are effectively propagated across different resolution levels, enhancing both global semantic consistency and fine-grained motion synchronization.
Finally, the denoised latent variable $Z_{0}$ is decoded into a sequence of video frames that not only exhibit dynamic facial expressions synchronized with the input audio, but also enable controllable facial attribute editing based on the selected semantic encoding $z_{\text{sem}}$, depending on the switch configuration. This process produces high-quality editable talking head videos. For further implementation details, please refer to Section IV-C.
III-E Training and Inference Pipeline
Training process: Our model employs a three-stage training framework to progressively learn semantic representation, attribute classification, and conditional generation. Each stage builds upon the previous to ensure robust feature learning, precise attribute control, high-quality and controllable synthesis.
The first stage serves as a pre-training phase, during which the semantic and stochastic encoders are jointly trained by minimizing the mean squared error between the predicted and ground-truth noise: